OpenStudy (anonymous):

describe the vertical asymptote(s) and hole(s) for the graph of y=x-4/x^2+3x+2

OpenStudy (anonymous):

the degree of the numerator is less than the degree of the denominator, so you have a horizontal asymptote at \(y=0\)

OpenStudy (anonymous):

i am assuming this is \[y=\frac{x-4}{x^2+3x+2}\]right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

to find the vertical asymptotes, set the denominator equal to zero and solve for \(x\( i.e. solve \[x^2+3x+2=0\]

OpenStudy (anonymous):

\[x^2+3x+2=0\]

OpenStudy (anonymous):

this one factors so it should be easy enough \[(x+2)(x+1)=0\] etc

OpenStudy (anonymous):

then you will set each of them equal to zero right?